OpenStudy (anonymous):

After a 5% pay raise, this years salary is 37,800 find last years salary?

OpenStudy (lynfran):

5/100*37800=??

OpenStudy (lynfran):

whats that equal too?

OpenStudy (lynfran):

then u take way that frm 37800

OpenStudy (anonymous):

or you could just subtract 5% from $37,800 & get the same answer

OpenStudy (anonymous):

I have a new problem now. I have been doing this same homework assignment for 3 hours and I'm so frustrated.

OpenStudy (anonymous):

This is the new problem

OpenStudy (zzr0ck3r):

\(x+0.05x=37800\) Solce for \(x\)

OpenStudy (lynfran):

@emily_wilson what answer did u get?

OpenStudy (anonymous):

for the first or second?

OpenStudy (lynfran):

the first one

OpenStudy (anonymous):

I got $35,910

OpenStudy (lynfran):

ok correct

OpenStudy (anonymous):

For the second one, I'm pretty sure you just take 3% & add it to $31.415

OpenStudy (lynfran):

the second 1 is the same process

OpenStudy (anonymous):

31,414.97?

OpenStudy (anonymous):

I got 32,357.45

OpenStudy (lynfran):

(3/100)*31415=942.45 then 31415-942.45=30.472.55

OpenStudy (anonymous):

I'm so done with this. I give up

OpenStudy (lynfran):

shd be 30,472.55

OpenStudy (anonymous):

it was 30,500

OpenStudy (lynfran):

yes when rounded

OpenStudy (anonymous):

I give up.

OpenStudy (anonymous):

why would you multiply on the second one?

OpenStudy (lynfran):

you find the amount for the 3% increase

OpenStudy (lynfran):

or is u used zzr0ck3r approach u would get that so x+0.03x=31415 1.03x=31415 x=31415/1.03 x=30500